OpenStudy (anonymous):

let g(x)=2x and h(x)=x-7 (g+h)(1/4)

OpenStudy (epoweritheta):

(g+h)x= g(x) +h(x) it s just a notation , now u do

OpenStudy (anonymous):

2x+x-7

OpenStudy (anonymous):

2x+x-7(1/4)

OpenStudy (epoweritheta):

yes...substitute x =1/4 in 2x+x-7

OpenStudy (anonymous):

2(1/4)+1/4-7

OpenStudy (epoweritheta):

super super simplify and tell

OpenStudy (anonymous):

0.5-6.75

OpenStudy (epoweritheta):

ya....

OpenStudy (epoweritheta):

next ??

OpenStudy (anonymous):

-6.25

OpenStudy (anonymous):

now im stuck

OpenStudy (epoweritheta):

thats the answer

OpenStudy (anonymous):

it says its wrong

OpenStudy (anonymous):

its right. thanks so much

OpenStudy (anonymous):

let g(x)=2x and h(x)=x-7 (g+h)(-1/5) same way for this one?

OpenStudy (epoweritheta):

ya same way it cant be wrong

OpenStudy (epoweritheta):

give me a medal pls and fan me for me to help u fast :)

OpenStudy (anonymous):

let g(x)=2x and h(x)=x-7 (g+h)(-1/5) g(x)+h(x) 2x+x-7(-1/5) 2(-1/5)+(-1/5)-7 answer : -7.6